Obtain the Fourier cosine series for the following function:

{1for 0≤ x<1
0 for 1≤ x <4 }
1
Expert's answer
2019-03-14T04:02:48-0400

Solution:

If

"f(x)=\begin{cases}
 1 &\text{ if } 0≤ x<1
 \\
 0 &\text{if 1≤ x <4 } 
\end{cases}"

then the Fourier cosine series for this function is:


"f(x)=\frac {a_0}{2}+{\displaystyle\sum_{n=1}^\infin}a_ncos\frac{n{\pi}x}{4},"




"0\u2264 x<4,"

where


"a_0=\frac{1}{2}{\smallint_0}^1 1dx+\frac{1}{2}{\smallint_1}^4 0dx=\frac{1}{2}"

and


"a_n=\frac{1}{2}{\smallint_0}^1 1cos\frac{n{\pi}x}{4}dx+\frac{1}{2}{\smallint_1}^4 0cos\frac{n{\pi}x}{4}dx=\frac{2}{n{\pi}}sin\frac{n{\pi}}{4}."

or


"f(x)=\frac {1}{4}+2{\displaystyle\sum_{n=1}^\infin}\frac{1}{n{\pi}}sin\frac{n{\pi}}{4}cos\frac{n{\pi}x}{4},""0\u2264 x<4."

Answer:

"f(x)=\frac {1}{4}+2{\displaystyle\sum_{n=1}^\infin}\frac{1}{n{\pi}}sin\frac{n{\pi}}{4}cos\frac{n{\pi}x}{4},"

"0\u2264 x<4."